For 2 consecutive integers, the result of adding the smaller integer and triple the larger integer is 79. What are the 2 integers?
Answer by checkley79(3341)

LET X & X+1 BE THE 2 INTEGERS.
X+3(X+1)=79
X+3X+3=79
4X=79-3
4X=76
X=76/4
X=19 ANS. FOR THE SMALLER INTEGER.
19+1=2- ANS. FOR THE LARGER INTEGER.
PROOF:
19+3*20=79
19+60=79
79=79